It takes a total of 475 seconds for one technician to assemble one AirRules smart phone, working alone. However, AirRules manufactures its smart phones with twenty technicians staffing a twenty station assembly line, producing one smart phone every 30 seconds. How efficient is the AirRules smart phone assembly line?

A. 50%

B. 21%

C. 33%

D. 67%

E. 79%
Answer #1

Time taken to assemble one smartphone = 475 secs

Time required by 20 technicians to assemble one phone = 475/20 = 23.75 secs

Actual time taken = 30 secs

Hence, efficiency = Time Required / Actual Time * 100% = 23.75/30 *100% = 79.17%

Hence, option (e) is correct option